OpenStudy (anonymous):

If x = 5, find -2times -6x

OpenStudy (phi):

Do you know how to do this?

OpenStudy (anonymous):

yes i just wanted to see if my answer was rite

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

thanks

OpenStudy (phi):

-2*-6*5= 12*5= 60

OpenStudy (kymber):

 Just put 5 in place of x and solve!